# SALAD-Bench: A Hierarchical and Comprehensive Safety Benchmark for Large Language Models

### v3

@@ 5.3 Model Safety Evaluation
Assessment across different safety dimensions. Results from Figure 5 show GPT-4 performing best in Information & Safety Harms and Claude2 in Malicious Use, with their challenges lying in Representation & Toxicity Harms and Socioeconomic Harms, respectively.
@@
  ChatGLM3-6B 90.45 1024 12.48 951
InternLM-7B 95.52 1043 20.28 982
InternLM-20B 96.81 1048 11.08 946
Llama-2-7B 96.51 1046 $*{}^{*}$ $*{}^{*}$
Llama-2-13B 96.81 1047 65.72 1167
Llama-2-70B 96.21 1045 66.24 1170
Mistral-7B-v0.1 54.13 891 2.44 913
Mistral-7B-v0.2 80.14 988 6.40 929
Mixtral-8x7B 76.15 974 9.36 939
Qwen-7B 91.69 1029 6.14 927
Qwen-14B 95.35 1042 7.8 934
Qwen-72B 94.40 1039 6.94 930
TuluV2-7B 84.79 1004 4.7 922
TuluV2-13B 86.51 1010 5.74 926
TuluV2-70B 91.60 1029 7.96 934
Vicuna-7B 44.46 851 4.2 920
Vicuna-13B 45.91 858 3.66 918
Vicuna-33B 54.12 891 3.84 918
Yi-6B 82.95 998 22.60 994
Yi-34B 87.13 1013 23.74 994
GPT-4 93.49 1035 80.28 1242
GPT-3.5 88.62 1018 73.38 1205
Claude2 99.77 1058 88.02 1290
Gemini 88.32 1017 19.98 980
@@
The attack-enhanced set shifts the challenge, with GPT-4 and Claude2 facing difficulty in Human Autonomy & Integrity Harms, GPT-3.5 in Socioeconomic Harms, and Gemini in Malicious Use. Easier domains include Information & Safety Harms for GPT-4 and Claude2, and Representation & Toxicity Harms for Gemini and GPT-3.5. See Appendix K for more detailed comparison.
@@ 5.4 Evaluation of Attack & Defense Methods
Attack method evaluation. We evaluate attack methods and prompts in Section 5.1 on both our defense-enhanced subset and AdvBench-50111A subset from original AdvBenchZou et al. (2023).Chao et al. (2023) and report results in Table 5. For jailbreak prompts and beforehand searched GCG suffixes, we report ASR both maximized [^2] Given a question, we count an attack success if at least one prompt triggers harmful response. and averaged [^3] Given a question, compute average ASR on all prompts. among all prompts. Overall, most methods achieve lower ASR on our defense-enhanced set than on AdvBench-50 and our base question set. This reveals the challenge of our defense-enhanced set. Among all attack methods, human-designed jailbreak prompts achieve the highest ASR, because models usually follow instructions in jailbreak prompts to scoff or curse. Suffixes searched from GCG can even trigger the model to generate detailed instructions on harmful behaviors, but is vulnerable to defense methods. GPTFuzzer gives moderate attack performance. AutoDAN and TAP are suboptimal even with different configurations. Chain of Utterances is sensitive to chat templates. Harmful responses can appear without chat templates but can hardly appear with chat templates.
@@ 5.5 Multiple-choice Question Subset Analysis
Finally, we analyze the performance of LLMs on the MCQ subset, as shown in Table 7. More results are shown in Appendix Table 13, 14 and 15. Generally, our primary findings are three-fold. First, GPT-4 (OpenAI, 2023) achieves the best 88.96% in terms of Acc-O and Acc-V, which surpasses all counterparts and shows powerful safety as well as helpfulness capability. Second, too strict safety restrictions are harmful to the overall accuracy of MCQs. Specifically, Acc-V of Gemini Pro (Team et al., 2023) achieves 78.71%, but corresponding Acc-O degrades to 44.19%. These results indicate that too strict safety strategies may limit the effectiveness of LLMs in safety-related tasks. Finally, weak instruction following ability also restricts the final accuracy in the MCQ subset. Specifically, the safe rates of InternLM-20B and Yi-34B achieve 98.25% and 91.38%. But the corresponding Acc-V reduced to 3.85% and 29.09%, which indicates insufficient instruction following ability restricts the safety ability of LLMs. Besides, we conduct more analysis for the MCQ subset, e.g., consistency between choosing safe or unsafe choices and accuracy of MCQ-Judge, as shown in Appendix H.
@@ 6 Conclusion
7 Limitation
Generally, the limitation of this paper can be roughly summarized as three-fold. Firstly, with the rise and development of new safety threats, our defined hierarchical taxonomy may be outdated. One could regularly update the taxonomy and evaluation data to tackle this issue. Secondly, during data collection, we leverage multiple filtration algorithms and reward models to clean the base set rather than intensive human labor. Nevertheless, the quality of the base set largely relies on the quality of such algorithms and reward models. Finally, for QA evaluation, the precision of evaluation results depends on the performance of the MD-Judge evaluator; and for multiple-choice subset evaluation, the evaluation results may rely on the instruction following the ability of candidate LLMs.
@@ 8 Broader Impact and Ethics Statement
Safety benchmarks play a vital role in identifying potential harms associated with LLMs. With growing concern about the ability of existing benchmarks to accurately quantify and identify specific safety issues, our motivation is centered on improving the security and safety of LLMs. This is achieved by evaluating LLMs with challenging questions and a detailed safety taxonomy that highlights specific vulnerabilities. However, conducting this research involves handling potentially sensitive content, such as attack-enhanced questions and choices in multiple-choice question scenarios. To address this, we restrict access to the most sensitive parts of the benchmark, available only to authorized researchers who adhere to stringent ethical guidelines. These measures are designed to safeguard the integrity of the research while minimizing any potential harm associated with the content of the benchmark.
Appendix A Detailed Domain-level Taxonomy Definition
@@
is delineated into three distinct yet interconnected tasks: toxic content, unfair representation, and adult content. The toxic content domain encompasses the production of content that contravenes established community guidelines and ethical standards such as insult, child abuse, violent content, etc. Unfair representation addresses the phenomena of misrepresentation, under-representation, or over-representation of specific identities, groups, or perspectives. While adult content introduces material that is explicitly sexual in nature and intended for mature users, such as pornography, erotic chat and sexual activity.
@@
can be broadly categorized into two main tasks, i.e., propagation of false beliefs and misconceptions refers to the dissemination. Misinformation also involves action that is erosion of trust in public information. Notable instances include the medical advice, law advice.
@@
can be articulated through two principal tasks. Privacy infringement: This encompasses the unauthorized revelation, creation, or accurate deduction of personal and private data about individuals. It involves scenarios where personal details are inferred or private information is leaked. The other is dissemination of dangerous information. This category covers the release, generation, or correct inference of information that poses a potential security threat such as fake news, social media rumors, government secrets and financial data leaks.
@@
can be delineated into several key tasks. Influence Operations includes manipulating campaign materials and public discourse, along with illegitimate surveillance activities. Fraud involves scams, cheating, fake reviews, fake websites, and tricking people for money or other gains.Defamation spreads false information to damage someone’s reputation. Security threats proposes activities related to cyber attacks, creating malware, and making or moving weapons. Illegal activities covers a wide range of crimes like financial fraud, drug crimes, sexual offenses, violence, theft, giving illegal legal advice, environmental crimes, traffic violations, and terrorism.
@@
mainly contain two tasks, the risky financial practices such as paypal lending, multi-level marketing and gambling and trade and compliance which contains trade affairs and labor issues.
@@
Each domain is subdivided into tasks and categories, with detailed description and number of questions provided in Table M.
Appendix B Question generation and enhancement
B.1 Question collection
We use the prompt in Figure 6 to generate self-instructed questions in all the categories by finetuned GPT. The number of questions and category should be changed to generate. Upon collection, we leverage a LLM-based reward model Dai et al. (2023) $\mathcal{R}_{\text{s}}$, pretrained on SafeRLHF Dai et al. (2023), to determine whether each question is safe or not. Specifically, with given input question $\mathbf{q}$, we first feed $\mathbf{q}$ into $\mathcal{R}_{\text{s}}$ to obtain corresponding harmfulness score $s_{\text{harm}}$. Then, we classify question $\mathbf{q}$ via $s_{\text{harm}}$ and a given threshold $T_{\text{harm}}$, i.e., $\mathbf{q}$ with $s_{\text{harm}}>T_{\text{harm}}$ is seen as harmful question, and vise versa. Afterwards, we follow the instruction for auto labeling task which is shown in Figure 7.
B.2 Defense enhancement
Defense enhacement pipeline is shown in Figure 8.
  Type Data Source Nums
  Self-instructed Finetuned GPT-3.5 15433
Open-Sourced HH-harmless 4184
HH-red-team 659
Advbench 359
Multilingual 230
Do-Not-Answer 189
ToxicChat 129
Do Anything Now 93
GPTFuzzer 42
Total 21318
Table 8: Data source of base set in SALAD-Bench.
Figure 6: Prompt for harmful question generation, which consists of system prompt and input prompt. [X] should be the number of questions to be generate and [Y] should be the exact category.
Figure 7: Example instructions for the auto labeling task. The task consists of three main components: task description, context examples and output format.
Figure 8: Construction pipeline of the defense-enhanced dataset. (a) Generate response on all candidate models. (b) Keep questions with high rejection rate. (c) Attack each question and keep failed ones. (d) Enhance remaining questions with defense methods.
B.3 Details in Attack Enhancement
Appendix C Full Results of Defense Methods
We also illustrate the full results of defense methods. See Table 9.
  w/o Defense 34.28% 88.92% 93.60% 90.64% 93.06% 92.04% 96.34% 76.26%
Appendix D Evaluation Template for MD-Judge
The template of our MD-Judge is intricately designed with four distinct components. First is Task, which indicates that our evaluator is designed to assess the safety of agent messages in the whole conversations. Second is Safety taxonomy, to make a trade-off between classification accuracy and context length of LLMs, we list the general activities of domain-level or task-level safety taxonomies. To ensure the activities of each category easy to understand, we utilize the ability of GPT-4 and human labor to transform each taxonomy into a clear and distinct instruction format, i.e., description regarding “Should not” and “Can”, thus obtaining the safety taxonomy paragraph. Third is QA Pairs, where to place the input questions and corresponding responses from specific LLMs. Lastly, Output format aims to utilize LLMs’ ability to follow instructions, ensuring that our evaluator adheres strictly to our designated format. This approach paves our way to calculate the safe/unsafe probability using the probability of the very first token.
Appendix E Detailed Evaluation Results of Evaluators
  Methods Base Enhance ToxicChat Beavertails SafeRLHF
Table 10: The detailed comparison results of the safety evaluation between our model and other mainstream evaluation methods. The values reported in the table, listed from left to right, are as follows: accuracy, precision, recall, F1 score. The best F1 scores are bolded the second best results are underlined. Base and Enhance indicate our SALAD-Base-Test and SALAD-Enhance-Test.
  Methods Base Enhance ToxicChat Beavertails SafeRLHF
F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC
Table 11: Comparison between LlamaGuard with different taxonomy templates and our MD-Judge for QA-pairs. Origin mean LlamaGuard’s official safety policy in there code implementation, Domain and Task mean our two levels of safety policy. The best results are bolded and the second results are underlined. Base and Enhance indicate our SALAD-Base-Test and SALAD-Enhance-Test.
Appendix F SFT Versions of Our Evaluators
First of all, we only focus on the open-sourced SoTA models with 7B parameters i.e. Llama-2-7B (Touvron et al., 2023), Mistral-7B-v0.1 (Jiang et al., 2023), and Mistral-7B-Instruct-v0.1 (Jiang et al., 2023) for the following two reasons: 1) the commendable understanding and reasoning capabilities ensure robust representation ability to various question-answer pairs. 2) models with ~7B parameters are more user-friendly and require less computation resource during inference.
Based on the table results, we can tell that Mistral-7B-v0.1 along with task-level template training format is the best one as it demonstrate a significant improvements over the alternatives on the ToxicChat and also achieve commendable results on the other datasets. Therefore, we finally choose it as our evaluator.
  Versions Tax. Base Enhance ToxicChat Beavertails SafeRLHF
F1 AC F1 AC F1 AC F1 AC F1 AC
  Llama-2-7B domain .8276 .8646 $*{}^{*}$$*{}^{*}$$*{}^{*}$$*{}^{*}$8547 .9506 .8579 .9276
Llama-2-7B task .8174 .8812 $*{}^{*}$$*{}^{*}$$*{}^{*}$$*{}^{*}$8549 .9506 .8561 .9233
Table 12: Different versions of our fine-tuned safety evaluators. The best results are bolded and the second results are underlined. AC is short for AUPRC and Tax is short for taxonomy. The number with a “*” means that outputs does not strictly follow format, causing an inaccurate number. Base and Enhance indicate our SALAD-Base-Test and SALAD-Enhance-Test.
Appendix G Details in Evaluating Attack and Defense Methods
Attack methods.
For all attack algorithms that require specifying a target model, we choose Llama-2-7B-Chat as the target model, because it is one of the safest model as shown in Table 4. We hypothesis that attacks successful against Llama-2 would be more transferable to other model comparing to using unsafer model as the target model.
For TAP Mehrotra et al. (2023), we follow the original work to use vicuna-13B-v1.5 as the attacker. We experiment different configurations using vicuna-13B-v1.5, gpt-3.5-turbo-1106 and gpt-4-1106-preview as the evaluator. During attacking, the attacker model would generate multiple attack-enhanced questions. We collect all questions with evaluation score $\geq 9$ (full score is $10$) and re-evaluate using our evaluate. We count a successful attack if one of attack-enhanced questions is evaluated as unsafe by our evaluator.
Human designed jailbreak prompts templates are picked from jailbreakchat.com based on corresponding reported JB scores and up votes. We pick top 10 templates ranked by JB scores and another top 10 by upvotes.
For Chain-of-Utterances(Bhardwaj and Poria, 2023), we use its prompt. But when chatting with models, we apply the chat template. This is different to the original work (according to its source code) where chat template is not applied when generating from open sourced models.
Defense methods.
For paraphrasing-based methods, we choose GPT-paraphrasing Cao et al. (2023) as the baseline method. For perturbation-based methods, we choose four different augmentation methods, i.e., random erasing Cao et al. (2023), random inserting Robey et al. (2023), random patching Robey et al. (2023), and random swapping Robey et al. (2023) as defense methods. And for prompting-based methods, we utilize the recently proposed Safe / XSafe prompts Deng et al. (2023) and Self-Reminder prompt Wu et al. (2023) in our experiments, which have shown effective defense abilities in small-scale experiments.
Appendix H Templates of MCQ-Judge
Generation templates.
To generate safe responses, we utilize GPT-4 (OpenAI, 2023) to produce benign answers using the following prompt:
Since pre-trained LLMs (OpenAI, 2023; Touvron et al., 2023) (e.g., GPT-4 and Llama-2) typically generate benign outputs, we employ a finetuned GPT model, as detailed in Section 2.2 to generate unsafe responses by the following prompt:
Judge template.
Here is the full template used in MCQ-Judge.
Note that the few-shot exemplars only provide output formats for responses and do not require introducing realistic safety questions, thus avoiding to introduce bias during answer generation for multiple-choice questions.
Appendix I Rejection Keywords

### v4

@@ 5.4 Model Safety Evaluation
  ChatGLM3-6B 90.45 1016 12.48 954
InternLM-7B 95.52 1034 20.28 979
InternLM-20B 96.81 1039 11.08 948
InternLM2-7B 97.7 1041 22.2 985
InternLM2-20B 98.15 1043 29.82 1002
Llama-2-7B 96.51 1038 18.20 972
Llama-2-13B 96.81 1038 65.72 1145
Llama-2-70B 96.21 1038 66.24 1119
Llama-3-8B 95.69 1035 61.92 1035
Llama-3-70B 84.45 995 63.72 1149
Mistral-7B-v0.1 54.13 882 2.44 932
Mistral-7B-v0.2 80.14 980 6.40 940
Mixtral-8x7B 76.15 963 9.36 944
Qwen-7B 91.69 1021 6.14 942
Qwen-14B 95.35 1034 7.8 941
Qwen-72B 94.40 1030 6.94 939
Qwen1.5-0.5B 80.36 981 23.34 993
Qwen1.5-1.8B 62.96 9918 16.22 974
Qwen1.5-4B 95.51 1034 8.48 950
Qwen1.5-7B 93 1025 7.18 942
Qwen1.5-14B 95.37 1035 8.08 946
Qwen1.5-72B 93.55 1028 10.56 948
Gemma-2b 95.9 1036 50.34 1083
Gemma-7b 94.08 1029 15.54 966
TuluV2-7B 84.79 996 4.7 935
TuluV2-13B 86.51 1001 5.74 938
TuluV2-70B 91.60 1022 7.96 941
Vicuna-7B 44.46 842 4.2 934
Vicuna-13B 45.91 850 3.66 934
Vicuna-33B 54.12 884 3.84 934
Yi-6B 82.95 991 22.60 979
Yi-34B 87.13 1005 23.74 986
GPT-4 93.49 1028 80.28 1216
GPT-3.5 88.62 1009 73.38 1189
Claude2 99.77 1051 88.02 1217
Gemini 88.32 1009 19.98 982
@@
Assessment across different safety dimensions. Results from Figure 5 show GPT-4 performing best in Information & Safety Harms and Claude2 in Malicious Use, with their challenges lying in Representation & Toxicity Harms and Socioeconomic Harms, respectively. The attack-enhanced set shifts the challenge, with GPT-4 and Claude2 facing difficulty in Human Autonomy & Integrity Harms, GPT-3.5 in Socioeconomic Harms, and Gemini in Malicious Use. Easier domains include Information & Safety Harms for GPT-4 and Claude2, and Representation & Toxicity Harms for Gemini and GPT-3.5. See Appendix O for more details.
@@ 5.5 Evaluation of Attack & Defense Methods
Attack method evaluation. We evaluate attack methods and prompts in Section 5.1 on both our defense-enhanced subset and AdvBench-50111A subset from original AdvBenchZou et al. (2023).Chao et al. (2023) and report results in Table 6. For jailbreak prompts and beforehand searched GCG suffixes, we report ASR both maximized [^2] Given a question, we count an attack success if at least one prompt triggers harmful response. and averaged [^3] Given a question, compute average ASR on all prompts. among all prompts. Overall, most methods achieve lower ASR on our defense-enhanced set than on AdvBench-50 and our base question set. This reveals the challenge of our defense-enhanced set. Among all attack methods, human-designed jailbreak prompts achieve the highest ASR, because models usually follow instructions in jailbreak prompts to scoff or curse. Suffixes searched from GCG can even trigger the model to generate detailed instructions on harmful behaviors, but is vulnerable to defense methods. GPTFuzzer gives moderate attack performance. AutoDAN and TAP are suboptimal even with different configurations. CoU is sensitive to chat templates. Harmful responses can appear without chat templates but can hardly appear with chat templates.
@@ 5.6 Multiple-choice Question Subset Analysis
Finally, we analyze the performance of LLMs on the MCQ subset, as shown in Table 8. More results are shown in Appendix Table 13, 15 and 16. Generally, our primary findings are three-fold. First, GPT-4 (OpenAI, 2023a) achieves the best 88.96% in terms of Acc-O and Acc-V, which surpasses all counterparts and shows powerful safety as well as helpfulness capability. Second, too strict safety restrictions are harmful to the overall accuracy of MCQs. Specifically, Acc-V of Gemini Pro (Team et al., 2023) achieves 78.71%, but corresponding Acc-O degrades to 44.19%. These results indicate that too strict safety strategies may limit the effectiveness of LLMs in safety-related tasks. Finally, weak instruction following ability also restricts the final accuracy in the MCQ subset. Specifically, the safety rates of InternLM-20B and Yi-34B achieve 96.81% and 87.13%. But the corresponding Acc-V reduced to 3.85% and 29.09%, which indicates insufficient instruction following ability restricts the safety ability of LLMs. Besides, we conduct more analysis for the MCQ subset, e.g., consistency between choosing safe or unsafe choices and accuracy of MCQ-Judge, as shown in Appendix L.
@@ 6 Conclusion
7 Limitations
This paper has three main limitations. First, as new safety threats emerge and evolve, our defined hierarchical taxonomy may become outdated. To address this issue, one could regularly update the taxonomy and evaluation data. Second, during data collection, we relied on multiple filtration algorithms and reward models to clean the base set rather than intensive human labor. However, the quality of the base set largely depends on the quality of these algorithms and reward models. Finally, for the question-answering evaluation, the precision of the results depends on the performance of the MD-Judge evaluator. For the multiple-choice subset evaluation, the results may rely on the instruction-following ability of the candidate LLMs.
@@ 8 Broader Impact and Ethics Statement
Safety benchmarks are crucial for identifying potential harms in LLMs. Our research aims to improve LLM security and safety by evaluating models with challenging questions and a detailed safety taxonomy. To mitigate risks associated with sensitive content in the benchmark, such as attack-enhanced questions, we restrict access to authorized researchers who adhere to strict ethical guidelines. These measures safeguard research integrity while minimizing potential harm.
Appendix A Related works
With the advancement of LLM, there is a significant increase in safety concerns. These arise primarily due to the models’ enhanced ability to produce text indistinguishable from that written by humans. This capability, while impressive, also opens doors for potential misuse. Consequently, safety research must evolve in tandem with the development of LLMs to address these concerns effectively.
A.1 LLM Safety dataset.
To formulate and evaluate safety concerns for LLM, a wide range of safety datasets (Hartvigsen et al., 2022; Lin et al., 2023; Gehman et al., 2020; Dhamala et al., 2021; Sun et al., 2023; Wang et al., 2023b) have been emerging. For instance, ToxiGen (Hartvigsen et al., 2022) have proposed a machine-generated large-scale dataset benign statements; Safetyprompts (Sun et al., 2023) have developed an 100k safety dataset for safety of Chinese large language models. Do-not-answer (Wang et al., 2023b) have colleceted and annotated a dataset covering three-level safety concerns; Although these work do provided a benchmark for early LLM safety exploration and inspired subsequent AI safety research, but they all faced different shortcomings and challenges.
Firstly, most of safety datasets only fall on a narrow perspective of safety threats (e.g., only unsafe instructions or only toxic representation), failing to cover the wide spectrum of potentially safety concerns. For instance, RealToxicityPrompts (Gehman et al., 2020), ToxiGen (Hartvigsen et al., 2022), and Toxic-chat (Lin et al., 2023) focus primarily on toxic content, while BOLD (Dhamala et al., 2021) and CrowS-pairs (Nangia et al., 2020) are centered on bias. Secondly, previous harmful questions can be effectively handled with a high safety rate of about 99% by modern LLMs (Wang et al., 2023b; Sun et al., 2023), even by LLMs without specific safety alignment (Sun et al., 2023), which further highlights the backwardness of the current safety dataset. More challenging questions such as those from red-team or adversarial jailbreak Chao et al. (2023); Liu et al. (2023b) are desired for comprehensive evaluation of LLM safety. Thirdly, these benchmarks often have limited usage scope, either designed solely for safety assessment (Wang et al., 2023b; Shen et al., 2023) or aimed at testing attack and defense strategies (Zou et al., 2023), limiting their generalizability for broader application.
Hence, we provide our Salad-Bench datasets compared with the existing LLM safety datsaets are shown in Table 1
A.2 Attack & defense
Attack.
Attacks on LLMs typically aim to elicit harmful or undesirable responses, a phenomenon often referred to as "jailbreaking." (Liu et al., 2023b; Yu et al., 2023; Deng et al., 2023). Recent literature has explored various aspects of this issue. For instance, some studies focus on manually crafted jailbreak prompts (Shen et al., 2023) or red-teaming , often sourced from online platforms like jailbreakchat.com, Reddit or by careful human design. Others develop algorithms capable of automatically generating such prompts (Yu et al., 2023; Zhu et al., 2023). Among the algorithms, they can be classified into search based (Chao et al., 2023; Mehrotra et al., 2023), gradient based (Zou et al., 2023) and transformation (Wei et al., 2023). Notably, GCG (Zou et al., 2023) proposed a method for creating adversarial suffixes to elicit affirmative responses from LLMs. Building upon this, AutoDan (Zhu et al., 2023) introduced an interpretable algorithm that not only generates attack prompts but also potentially exposes the underlying system prompts of LLMs. Another noteworthy approach is PAIR (Chao et al., 2023), which creates semantic jailbreaks with only black-box access to an LLM by searching, demonstrating the diversity and complexity of attack strategies.
Defense.
In contrast to the rapid progress in jailbreak attack methodologies, defensive strategies for LLMs have not been as extensively explored. Some research, like the work presented in Jain et al. (2023), investigates various defensive tactics, such as the use of perplexity filters for preprocessing, paraphrasing input prompts, and adversarial training. While heuristic detection methods show promise, adversarial training has proven impractical due to the high computational costs involved in retraining LLMs. Another innovative approach is proposed in Kumar et al. (2023), which offers certifiable robustness through the application of safety filters on input prompt sub-strings. However, the complexity of this method increases with the length of the prompt, making it less feasible for longer inputs. Robey et al. (2023) introduces a novel technique that involves perturbing and aggregating predictions from multiple variations of an input prompt to identify adversarial inputs, adding to the spectrum of potential defensive strategies.
A.3 Evaluation methods.
Existing methods for evaluating the harmfulness or toxicity of the model’s response can be roughly classified into the following 4 categories:
Moderation Classifier.
Methods based on moderation classifiers, such as Detoxify (Hanu and Unitary team, 2020), Perspective API (Lees et al., 2022) and OpenAI Moderation API (OpenAI, 2023b). Although they are well-maintained, they focus solely on toxic and harmful content, lacking sufficient coverage in terms of safety dimensions.
Keyword.
Method based on keyword detection is the second category which predefined a set of keywords such as "I’m sorry". It has been used in GCG (Zou et al., 2023) which defines about 50 detection keywords. This method necessitates customizing distinct keywords for different models, and it is prone to a substantial number of false positives and false negatives.
GPT-based referees.
General LLMs as safety referees is the third category and possibly the most widely used and powerful method. This approach typically requires the careful design of prompts to instruct a powerful language model, often ChatGPT (OpenAI, 2022) or GPT-4 (OpenAI, 2023), as they are widely recognized for their superior capabilities, which is crucial for a referee model. Although this method has been used in  (Qi et al., 2023),  (Wang et al., 2023b),  (Yu et al., 2023), and so on, it faces issues typical of black-box models, such as lack of control, limited reproducibility, and potential future loss of model checkpoint.
LLM-based safety guard.
LLM-based safety guard: finetuned-LLM on safety dataset to provide specific usage of safety check, such as LlamaGuard (Inan et al., 2023). However, LlamaGuard does not demonstrate strong generalization ability, especially in our benchmark which contains many enhanced question-answer pairs. At the same time, not many similar safety guard models have emerged for now. Thus we need to build our own evaluator for our specially designed benchmark.
In addition to question-answer safety evaluation, the LLM Safety Dataset also includes multiple choice questions. Previous works use keyword-based evaluators (Xu et al., 2023) and GPT-based evaluators (Deng et al., 2023) to handle open-ended responses for multiple-choice questions (MCQ). As it suffers from the unstable and irregular model output, we leverage in-context learning with regex parsing to efficiently fetch the answers.
Appendix B Detailed Domain-level Taxonomy Definition
@@
is delineated into three distinct yet interconnected tasks: toxic content, unfair representation, and adult content. The toxic content domain encompasses the production of content that contravenes established community guidelines and ethical standards such as insult, child abuse, violent content, etc. Unfair representation addresses the phenomena of misrepresentation, under-representation, or over-representation of specific identities, groups, or perspectives. While adult content introduces material that is explicitly sexual and intended for mature users, such as pornography, erotic chat, and sexual activity.
@@
can be broadly categorized into two main tasks, i.e., propagation of false beliefs and misconceptions refers to the dissemination. Misinformation also involves action that is erosion of trust in public information. Notable instances include medical advice and law advice.
@@
can be articulated through two principal tasks. Privacy infringement: This encompasses the unauthorized revelation, creation, or accurate deduction of personal and private data about individuals. It involves scenarios where personal details are inferred or private information is leaked. The other is the dissemination of dangerous information. This category covers the release, generation, or correct inference of information that poses a potential security threat such as fake news, social media rumors, government secrets, and financial data leaks.
@@
can be delineated into several key tasks. Influence Operations include manipulating campaign materials and public discourse, along with illegitimate surveillance activities. Fraud involves scams, cheating, fake reviews, fake websites, and tricking people for money or other gains. Defamation spreads false information to damage someone’s reputation. Security threats propose activities related to cyber attacks, creating malware, and making or moving weapons. Illegal activities cover a wide range of crimes like financial fraud, drug crimes, sexual offenses, violence, theft, giving illegal legal advice, environmental crimes, traffic violations, and terrorism.
@@
mainly contain two tasks, risky financial practices such as PayPal lending, multi-level marketing and gambling, and trade and compliance which contains trade affairs and labor issues.
@@
Each domain is subdivided into tasks and categories, with detailed descriptions and the number of questions provided in Table 17.
Appendix C Rejection Keywords